A small company produces both standard and deluxe playhouses. The standard playhouses take 12 hours of labor to produce, and the deluxe playhouses take 20 hours. The labor available is limited to 800 hours per week, and the total production capacity is 50 items per week. Existing orders require the company to produce at least 10 standard playhouses and 15 deluxe playhouses per week. Write a system of inequalities representing this situation, where x is the number of standard playhouses and y is the number of deluxe playhouses. Then graph the system of inequalities
Answer by scott8148(6628)

an x takes 12 hours to build and a y takes 20 hours ... there are 800 hours available ... 12x+20y<=800 or 3x-5y<=200
total capacity is 50 ... x+y<=50
to cover exsisting orders ... x>=10, y>=15
graph the inequalities like equations ... after you draw the line, pick a point on one side and see if it satisfies the inequality - if it does, then shade that side as the solution - if not, shade the other side